OpenStudy (anonymous):

how do you do 12x^2-8x-15=0?

OpenStudy (anonymous):

Look for two numbers that multiply to -180 and add to -8 (-18 and 10). then factor by grouping and finally use the zero factor property.

OpenStudy (anonymous):

i got (x^2+10) (-18-180) =0

OpenStudy (anonymous):

Close just remember that when you rewrite it, it is x^2+10x-18x-180 and then factor in groups

OpenStudy (anonymous):

(12x^2+10x)+(-18x-180) You have this first, now can you factor out the like terms in each set of brackets?

OpenStudy (anonymous):

3x? is the like term?

OpenStudy (anonymous):

no, for the first bracket 2x will be the like term because 2x goes into 12x^2 and 10x

OpenStudy (anonymous):

Sorry I have a typo it should be -15 not 180

OpenStudy (anonymous):

You want to split the middle term of 12x^2-8x-15 as 12x^2-18x+10x-15

OpenStudy (anonymous):

for the second brackets the common factor would be 3 right?

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

Correct -3 :)

OpenStudy (anonymous):

so, is that the answer? -3 and 2? Or i have to do more? Because of the 0?

OpenStudy (anonymous):

So far we have this 2x(6x+5)-(6x+5)=0 We can factor the 6x+5 to get (6x+5)(2x-1)=0 Then we can solve for x

OpenStudy (anonymous):

What answers do you get?

OpenStudy (anonymous):

x=5/6 and x=-1/2???

OpenStudy (anonymous):

that's not right is it?

OpenStudy (anonymous):

No, right numbers just the wrong sign. for the 2x-1 bracket I have 2x-1=0 2x=1 x=1/2 Do you see this? 6x+5=0 6x=-5 x=-5/6

OpenStudy (anonymous):

Really close though, good work

OpenStudy (anonymous):

so i basically got it right, i just had to switch the negative and pos. signs? Thanks for helping me!

OpenStudy (anonymous):

No worries :)

OpenStudy (anonymous):

do you think you could help me with x^2-29x+204?

OpenStudy (anonymous):

sure I posted on that question lets move to it now